# NeurIPS 2025 Workshop on Space in Vision, Language, and Embodied AI

### v1

@@ Abstract
Common computer vision systems typically assume ideal pinhole cameras but fail when facing real-world camera effects such as fisheye distortion and rolling shutter, mainly due to the lack of learning from training data with camera effects. Existing data generation approaches suffer from either high costs, sim-to-real gaps or fail to accurately model camera effects. To address this bottleneck, we propose 4D Gaussian Ray Tracing (4D-GRT), a novel two-stage pipeline that combines 4D Gaussian Splatting with physically-based ray tracing for camera effect simulation. Given multi-view videos, 4D-GRT first reconstructs dynamic scenes, then applies ray tracing to generate videos with controllable, physically accurate camera effects. 4D-GRT achieves the fastest rendering speed while performing better or comparable rendering quality compared to existing baselines. Additionally, we construct eight synthetic dynamic scenes in indoor environments across four camera effects as a benchmark to evaluate generated videos with camera effects.
@@ 1 Introduction
However, common computer vision systems only consider pinhole cameras, rather than leveraging the beneficial properties of camera effects, because of the lack of high-quality visual data paired with accurate camera-effect parameters. As a result, these systems could fail when tested with these common and unavoidable camera effects in real-world scenarios. Gong et al. (2024); Fang et al. (2024); Fang and Chen (2025); Wang et al. (2023b) This data scarcity issue becomes a significant obstacle for modeling these effects, particularly in dynamic scenes where they interact non-trivially with motion (e.g., rolling shutter depends on object and scene dynamics Zhong et al. (2021); Saurer et al. (2013)). To overcome this issue, a straightforward approach is to synthesize data with camera effects. Traditional methods rely on physics-based rendering engines such as Blender Blender Online Community (2025), but they are labor intensive, costly, and can suffer from a pronounced sim-to-real gap Nikolenko et al. (2021); Tremblay et al. (2018). More recently, world models that are largely built on pretrained video generation models Ball et al. (2025); He et al. (2025); Hu et al. (2023, 2024) have been promoted as a path to realistic, controllable data. Unfortunately, as shown in Fig. 2, off-the-shelf video generation models have limited understanding of camera effects and their corresponding parameters, often generating videos that violate physical principles. These findings suggest that such world models underfit visual data paired with accurate camera-effect parameters, creating a chicken-and-egg problem.
The failure of these approaches motivates us to generate such video data by direct 4D reconstruction. With a reconstructed 4D representation, we can eliminate both the high cost of real-world data collection and the sim-to-real gap. However, existing reconstruction-based methods present significant shortcomings for this task. While Dynamic Gaussian Splatting methods Luiten et al. (2024); Yang et al. (2024b); Wu et al. (2024a) excel at dynamic scene reconstruction, they lack the capability to simulate light transport necessary for camera effect modeling. Conversely, Dynamic NeRF approaches Pumarola et al. (2021); Park et al. (2021a); Gao et al. (2021); Park et al. (2021b); Fridovich-Keil et al. (2023); Fang et al. (2022b); Cao and Johnson (2023) can model light transport but suffer from prohibitively slow rendering speed, along with inferior reconstruction quality compared to Gaussian Splatting methods Kerbl et al. (2023).
@@ 2.1 Generating Visual Data with Camera Effect
Synthetic Data Generation Approaches. Traditional approaches for generating camera effect data rely heavily on synthetic rendering pipelines using computer graphics software such as Blender Blender Online Community (2025). These methods provide precise control over camera parameters, generating numerically labeled ground truth data. However, such methods face several fundamental limitations: (1) high labor costs for creating realistic scenes, (2) a sim-to-real gap that limits model generalization.
Generative Model Approaches. The emergence of powerful generative models, particularly diffusion-based methods Rombach et al. (2022); Croitoru et al. (2023); Blattmann et al. (2023); Xing et al. (2024); Ho et al. (2022a); Chang et al. (2025); Ho et al. (2022b); Chu et al. (2023); Yeh et al. (2024), has opened new avenues for camera effect synthesis Fang et al. (2024). These approaches can generate diverse visual content with various camera effects. Recent representative works include Curved Diffusion Voynov et al. (2024), AKiRa Wang et al. (2025a), which enable control over optical distortions or camera effect parameters. However, generative methods suffer from several critical limitations for training data generation: (1) most focus on specific camera effects rather than general effect synthesis, (2) lack of precise physical constraints leading to unrealistic camera effects, (3) inability to accurately control camera parameters, and (4) difficulty in ensuring multi-view consistency or temporal consistency in dynamic scenes.
@@ 2.2 Dynamic Neural Rendering
NeRF. Neural Radiance Fields (NeRF) Mildenhall et al. (2021) introduced a new paradigm for representing static scenes, modeling them as continuous volumetric functions with MLPs that map 3D coordinates and viewing directions to density and color for photorealistic view synthesis. Numerous extensions Barron et al. (2021); Zhang et al. (2020); Yu et al. (2021); Müller et al. (2022); Sun et al. (2022); Barron et al. (2022); Su et al. (2024); Lin et al. (2025) address NeRF’s limitations, including adaptations for dynamic scenes Pumarola et al. (2021); Park et al. (2021a, b); Gao et al. (2021); Fang et al. (2022a); Liu et al. (2023); Yan et al. (2023); Song et al. (2023); Gan et al. (2023); Cao and Johnson (2023); Wang et al. (2023a); Chen et al. (2024). Among these, K-Planes Fridovich-Keil et al. (2023), HexPlane Cao and Johnson (2023), and MSTH Wang et al. (2023a) achieve state-of-the-art performance by leveraging plane- or grid-based representations to accelerate dynamic scene reconstruction while maintaining high rendering quality.
Gaussian Splatting. Despite NeRF’s impressive results, its high computational cost and implicit representation limit practical use. 3D Gaussian Splatting (3D-GS)Kerbl et al. (2023) addresses these issues with an explicit 3D Gaussian Ke et al. (2025) representation and efficient rasterization, enabling real-time, high-quality rendering Hou et al. (2025); Zhan et al. (2025). Recent worksYang et al. (2024b); Wu et al. (2024b); Liang et al. (2025); Huang et al. (2023); Lin et al. (2024); Kratimenos et al. (2024); Yang et al. (2024a); Fan et al. (2024) extend 3D-GS to dynamic scenes. For example, Deformable 3D Gaussians Yang et al. (2024b) introduce deformation modeling, while 4D Gaussian Splatting (4D-GS) Wu et al. (2024b) employs a spatial-temporal encoder composed of multi-resolution voxel planes and MLPs to model the deformation of 3D Gaussians. These advances preserve the efficiency of Gaussian Splatting while enabling high-quality dynamic scene reconstruction, establishing a strong foundation for camera effect simulation.
@@ 2.3 Camera effect rendering in neural scene representations
Traditional renderers simulate camera effects through ray tracing Cook et al. (1984), reproducing physical optics. In neural scene representations, NeRF can achieve similar effects Wu et al. (2022); Niu et al. (2024) due to its ray-casting–based rendering, but NeRF-based methods are computationally expensive and lack efficient rendering. In contrast, 3D-GS offers fast rasterization but is restricted to the pinhole camera model, requiring specialized designs and complex Jacobian computations to support effects such as fisheye distortion Liao et al. (2024), rolling shutter with motion blur Seiskari et al. (2024), and depth of field Lee et al. (2025); Wang et al. (2025b); Shen et al. (2025). 3DGRT Moenne-Loccoz et al. (2024b) overcomes this limitation by enabling direct ray tracing over 3D Gaussians, allowing physically accurate and flexible camera effect simulation, while 3DGUT Wu et al. (2025) achieves similar capabilities by approximating 3D Gaussians with sigma points via the unscented transform and projecting them through different camera models. However, both are restricted to static scenes and cannot simulate camera effects in dynamic settings. To address this, our method integrates 4D-GS with 3DGRT, leveraging ray tracing to enable camera effect rendering in dynamic scenes.
@@ 3 Method
As illustrated in Fig.3, our proposed 4D Gaussian Ray Tracing (4D-GRT) framework generates physically correct training data in two stages: dynamic scene reconstruction (Section 3.2) and camera-effect rendering (Section 3.3). In the first stage, we reconstruct a dynamic scene from synchronized multi-view videos using a 4D-GS Wu et al. (2024a) model with differentiable ray tracing Moenne-Loccoz et al. (2024a). In the second stage, we render the scene with physically based camera models, including fisheye distortion, depth of field, and rolling shutter, producing physically accurate videos for effect-aware vision training.
@@ 3.1 Preliminary
3D Gaussian Splatting (3D-GS). 3D-GS Kerbl et al. (2023) represents a static 3D scene as a collection of 3D Gaussians, each characterized by its mean position $(x,y,z)$, a covariance matrix $\mathbf{\Sigma}$ decomposed into scaling $s$ and rotation $r$ parameters, opacity $\sigma$, and spherical harmonic coefficients $\mathcal{C}$. Rendering is done by differentiable splatting Yifan et al. (2019), where $\mathbf{\Sigma}$ is projected into 2D as $\mathbf{\Sigma}^{\prime}=\mathbf{J}\mathbf{W}\mathbf{\Sigma}\mathbf{W}^{T}\mathbf{J}^{T}$, with $\mathbf{J}$ and $\mathbf{W}$ denoting the Jacobian of the projective transformation and the viewing transformation, respectively. Pixel colors are then computed via alpha compositing of the projected Gaussians from front to back, $C=\sum_{i\in\mathcal{N}}T_{i}\alpha_{i}c_{i}$, where $T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j})$ is the transmittance, $\alpha_{i}=\sigma_{i}\exp(-\tfrac{1}{2}\bm{x}^{T}\mathbf{\Sigma}^{\prime}\bm{x})$ is the opacity, and $c_{i}$ is the view-dependent color derived from spherical harmonics.
@@ 3.2 Dynamic Scene Reconstruction
As shown in Fig.  3, we reconstruct scene dynamics from multi-view videos $\{\hat{C}_{v,t}\mid v\in\mathcal{V},t\in\mathcal{T}\}$, where $\mathcal{V}$ denotes the set of camera views and $\mathcal{T}$ the set of time indices. Following 4D-GS Wu et al. (2024a), we represent the scene with a set of 3D Gaussians $G$ and a Gaussian deformation field network, which consists of a spatio-temporal structure encoder and a multi-head Gaussian deformation decoder to model the per-frame deformations of 3D Gaussians. Specifically, to compute the deformation of 3D Gaussians $G$ at time $t$, we first employ a 4D voxel planes $R_{l}$ to extract voxel features based on each Gaussian’s mean position and time step $t$. These features are first fused by a lightweight MLP $\varphi$, and then passed to separate MLPs $(\varphi_{x},\varphi_{r},\varphi_{s})$ that predict the residuals of the Gaussian attributes $((\Delta x,\Delta y,\Delta z),\Delta r,\Delta s)$. Finally, the attribute residuals are applied to $G$ to obtain the deformed Gaussians $G_{t}$. For details of the Gaussian deformation field network, please refer to the supplementary material.
For rendering, although 3D-GS Kerbl et al. (2023) rasterization provides efficient and high-quality results, it is inherently limited in its ability to simulate physically accurate camera effects. In particular, phenomena that require explicit ray-based modeling, such as complex lens distortions or light-transport-dependent optical effects, cannot be faithfully reproduced. To address these limitations, we adopt the differentiable ray tracing framework of 3DGRT Moenne-Loccoz et al. (2024a), which traces rays directly through 3D Gaussian primitives instead of projecting them onto the image plane. By combining a $k$-buffer hit-based marching scheme with the hardware-accelerated NVIDIA OptiX interface, this approach achieves both the accuracy required for realistic camera-effect rendering and the efficiency necessary for large-scale data generation.
By integrating the dynamic scene model with differentiable ray tracing, we reconstruct dynamic scenes end-to-end via training with a color loss. At each training iteration, we randomly sample a view $v\in\mathcal{V}$ and a time $t\in\mathcal{T}$, obtain the deformed Gaussians $G_{t}$, and render them via ray tracing from view $v$ to obtain the image $C_{v,t}$. We then compute the L1 loss between $C_{v,t}$ and the ground-truth image $\hat{C}_{v,t}$, denoted by $\mathcal{L}_{1}$. Additionally, we apply a grid-based total variation loss $\mathcal{L}_{\mathrm{TV}}$ following Wu et al. (2024a); Cao and Johnson (2023). The overall loss function is defined as
@@ 3.3 Camera Effects Rendering
Fisheye. Fisheye lenses enable extremely wide-angle capture but introduce strong radial distortion. To simulate this effect, we follow Blender Blender Documentation Team (2025) and adopt a 4th-degree polynomial radial distortion model, which provides a flexible and physically grounded representation of fisheye lens characteristics. The model is defined as:
@@
Depth of Field (DoF). DoF controls which parts of a scene appear sharp and which appear blurred, allowing cameras to emphasize subjects at specific depths while de-emphasizing background or foreground regions. We simulate this effect using the model introduced in Cook et al. (1984). For each pixel, given the focus distance $f_{z}$ and the aperture radius $r_{a}$, we first compute the intersection point $\mathbf{p}=\mathbf{o}+f_{z}\mathbf{d}$ of the ideal pinhole ray with the focal plane at distance $f_{z}$. We then jitter the ray origin over a circular aperture of radius $r_{a}$ by sampling a point $\bm{\ell}=(\ell_{x},\ell_{y})$ uniformly from the unit disk and mapping it to world space using the camera’s lens-plane basis $(\mathbf{u},\mathbf{v})$, where $\mathbf{u}$ and $\mathbf{v}$ are orthogonal unit vectors spanning the lens plane. The perturbed ray origin and direction are then computed as
@@
Rolling Shutter. The rolling shutter effect arises because image rows are not captured simultaneously but sequentially over time. As a result, objects in motion or a moving camera can cause geometric distortions in the recorded image. To simulate this effect, we cast a single ray for each pixel through the deformed 3D Gaussians at its corresponding capture time. This approach resembles the motion blur technique of Cook et al. (1984), but it differs in that we assume no motion blur: instead of integrating multiple rays over the shutter interval of a row, we trace only one ray per pixel at the corresponding sensing time. Concretely, for a pixel at row $r$, we first compute the sensing time $t_{r}$, then evaluate the deformed 3D Gaussians $G_{t_{r}}$. A ray is traced through the deformed 3D Gaussians $G_{t_{r}}$ to determine the pixel’s radiance. We currently assume a static camera; however, the method can be extended to a moving camera by interpolating the camera trajectory at time $t_{r}$ and casting the ray from the position.

### v2

@@ Abstract
Common computer vision systems typically assume ideal pinhole cameras but fail when facing real-world camera effects such as fisheye distortion and rolling shutter, mainly due to the lack of learning from training data with camera effects. Existing data generation approaches suffer from either high costs, sim-to-real gaps or fail to accurately model camera effects. To address this bottleneck, we propose 4D Gaussian Ray Tracing (4D-GRT), a novel two-stage pipeline that combines 4D Gaussian Splatting with physically-based ray tracing for camera effect simulation. Given multi-view videos, 4D-GRT first reconstructs dynamic scenes, then applies ray tracing to generate videos with controllable, physically accurate camera effects. 4D-GRT achieves the fastest rendering speed while performing better or comparable rendering quality compared to existing baselines. Additionally, we construct eight synthetic dynamic scenes in indoor environments across four camera effects as a benchmark to evaluate generated videos with camera effects. Project page: https://shigon255.github.io/4DGRT-project-page.
@@ 1 Introduction
However, common computer vision systems only consider pinhole cameras, rather than leveraging the beneficial properties of camera effects, because of the lack of high-quality visual data paired with accurate camera-effect parameters. As a result, these systems could fail when tested with these common and unavoidable camera effects in real-world scenarios. [27, 19, 18, 73] This data scarcity issue becomes a significant obstacle for modeling these effects, particularly in dynamic scenes where they interact non-trivially with motion (e.g., rolling shutter depends on object and scene dynamics [91, 64]). To overcome this issue, a straightforward approach is to synthesize data with camera effects. Traditional methods rely on physics-based rendering engines such as Blender [7], but they are labor intensive, costly, and can suffer from a pronounced sim-to-real gap [55, 70]. More recently, world models that are largely built on pretrained video generation models [2, 30, 35, 36] have been promoted as a path to realistic, controllable data. Unfortunately, as shown in Fig. 2, off-the-shelf video generation models have limited understanding of camera effects and their corresponding parameters, often generating videos that violate physical principles. These findings suggest that such world models underfit visual data paired with accurate camera-effect parameters, creating a chicken-and-egg problem.
The failure of these approaches motivates us to generate such video data by direct 4D reconstruction. With a reconstructed 4D representation, we can eliminate both the high cost of real-world data collection and the sim-to-real gap. However, existing reconstruction-based methods present significant shortcomings for this task. While Dynamic Gaussian Splatting methods [50, 84, 77] excel at dynamic scene reconstruction, they lack the capability to simulate light transport necessary for camera effect modeling. Conversely, Dynamic NeRF approaches [60, 58, 25, 59, 23, 21, 10] can model light transport but suffer from prohibitively slow rendering speed, along with inferior reconstruction quality compared to Gaussian Splatting methods [40].
@@ 2.1 Generating Visual Data with Camera Effect
Synthetic Data Generation Approaches. Traditional approaches for generating camera effect data rely heavily on synthetic rendering pipelines using computer graphics software such as Blender [7]. These methods provide precise control over camera parameters, generating numerically labeled ground truth data. However, such methods face several fundamental limitations: (1) high labor costs for creating realistic scenes, (2) a sim-to-real gap that limits model generalization.
Generative Model Approaches. The emergence of powerful generative models, particularly diffusion-based methods [62, 15, 5, 81, 32, 11, 33, 13, 85], has opened new avenues for camera effect synthesis [19]. These approaches can generate diverse visual content with various camera effects. Recent representative works include Curved Diffusion [71], AKiRa [74], which enable control over optical distortions or camera effect parameters. However, generative methods suffer from several critical limitations for training data generation: (1) most focus on specific camera effects rather than general effect synthesis, (2) lack of precise physical constraints leading to unrealistic camera effects, (3) inability to accurately control camera parameters, and (4) difficulty in ensuring multi-view consistency or temporal consistency in dynamic scenes.
@@ 2.2 Dynamic Neural Rendering
NeRF. Neural Radiance Fields (NeRF) [51] introduced a new paradigm for representing static scenes, modeling them as continuous volumetric functions with MLPs that map 3D coordinates and viewing directions to density and color for photorealistic view synthesis. Numerous extensions [3, 89, 87, 54, 69, 4, 68, 47] address NeRF’s limitations, including adaptations for dynamic scenes [60, 58, 59, 25, 20, 49, 82, 67, 24, 10, 72, 12]. Among these, K-Planes [23], HexPlane [10], and MSTH [72] achieve state-of-the-art performance by leveraging plane- or grid-based representations to accelerate dynamic scene reconstruction while maintaining high rendering quality.
Gaussian Splatting. Despite NeRF’s impressive results, its high computational cost and implicit representation limit practical use. 3D Gaussian Splatting (3D-GS)[40] addresses these issues with an explicit 3D Gaussian [39] representation and efficient rasterization, enabling real-time, high-quality rendering [34, 88]. Recent works[84, 78, 45, 37, 48, 42, 83, 17] extend 3D-GS to dynamic scenes. For example, Deformable 3D Gaussians [84] introduce deformation modeling, while 4D Gaussian Splatting (4D-GS) [78] employs a spatial-temporal encoder composed of multi-resolution voxel planes and MLPs to model the deformation of 3D Gaussians. These advances preserve the efficiency of Gaussian Splatting while enabling high-quality dynamic scene reconstruction, establishing a strong foundation for camera effect simulation.
@@ 2.3 Camera effect rendering in neural scene representations
Traditional renderers simulate camera effects through ray tracing [14], reproducing physical optics. In neural scene representations, NeRF can achieve similar effects [80, 56] due to its ray-casting–based rendering, but NeRF-based methods are computationally expensive and lack efficient rendering. In contrast, 3D-GS offers fast rasterization but is restricted to the pinhole camera model, requiring specialized designs and complex Jacobian computations to support effects such as fisheye distortion [46], rolling shutter with motion blur [65], and depth of field [43, 75, 66]. 3DGRT [53] overcomes this limitation by enabling direct ray tracing over 3D Gaussians, allowing physically accurate and flexible camera effect simulation, while 3DGUT [79] achieves similar capabilities by approximating 3D Gaussians with sigma points via the unscented transform and projecting them through different camera models. However, both are restricted to static scenes and cannot simulate camera effects in dynamic settings. To address this, our method integrates 4D-GS with 3DGRT, leveraging ray tracing to enable camera effect rendering in dynamic scenes.
@@ 3 Method
As illustrated in Fig.3, our proposed 4D Gaussian Ray Tracing (4D-GRT) framework generates physically correct training data in two stages: dynamic scene reconstruction (Section 3.2) and camera-effect rendering (Section 3.3). In the first stage, we reconstruct a dynamic scene from synchronized multi-view videos using a 4D-GS [77] model with differentiable ray tracing [52]. In the second stage, we render the scene with physically based camera models, including fisheye distortion, depth of field, and rolling shutter, producing physically accurate videos for effect-aware vision training.
@@ 3.1 Preliminary
3D Gaussian Splatting (3D-GS). 3D-GS [40] represents a static 3D scene as a collection of 3D Gaussians, each characterized by its mean position $(x,y,z)$, a covariance matrix $\mathbf{\Sigma}$ decomposed into scaling $s$ and rotation $r$ parameters, opacity $\sigma$, and spherical harmonic coefficients $\mathcal{C}$. Rendering is done by differentiable splatting [86], where $\mathbf{\Sigma}$ is projected into 2D as $\mathbf{\Sigma}^{\prime}=\mathbf{J}\mathbf{W}\mathbf{\Sigma}\mathbf{W}^{T}\mathbf{J}^{T}$, with $\mathbf{J}$ and $\mathbf{W}$ denoting the Jacobian of the projective transformation and the viewing transformation, respectively. Pixel colors are then computed via alpha compositing of the projected Gaussians from front to back, $C=\sum_{i\in\mathcal{N}}T_{i}\alpha_{i}c_{i}$, where $T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j})$ is the transmittance, $\alpha_{i}=\sigma_{i}\exp(-\tfrac{1}{2}\bm{x}^{T}\mathbf{\Sigma}^{\prime}\bm{x})$ is the opacity, and $c_{i}$ is the view-dependent color derived from spherical harmonics.
@@ 3.2 Dynamic Scene Reconstruction
As shown in Fig.  3, we reconstruct scene dynamics from multi-view videos $\{\hat{C}_{v,t}\mid v\in\mathcal{V},t\in\mathcal{T}\}$, where $\mathcal{V}$ denotes the set of camera views and $\mathcal{T}$ the set of time indices. Following 4D-GS [77], we represent the scene with a set of 3D Gaussians $G$ and a Gaussian deformation field network, which consists of a spatio-temporal structure encoder and a multi-head Gaussian deformation decoder to model the per-frame deformations of 3D Gaussians. Specifically, to compute the deformation of 3D Gaussians $G$ at time $t$, we first employ a 4D voxel planes $R_{l}$ to extract voxel features based on each Gaussian’s mean position and time step $t$. These features are first fused by a lightweight MLP $\varphi$, and then passed to separate MLPs $(\varphi_{x},\varphi_{r},\varphi_{s})$ that predict the residuals of the Gaussian attributes $((\Delta x,\Delta y,\Delta z),\Delta r,\Delta s)$. Finally, the attribute residuals are applied to $G$ to obtain the deformed Gaussians $G_{t}$. For details of the Gaussian deformation field network, please refer to the supplementary material.
For rendering, although 3D-GS [40] rasterization provides efficient and high-quality results, it is inherently limited in its ability to simulate physically accurate camera effects. In particular, phenomena that require explicit ray-based modeling, such as complex lens distortions or light-transport-dependent optical effects, cannot be faithfully reproduced. To address these limitations, we adopt the differentiable ray tracing framework of 3DGRT [52], which traces rays directly through 3D Gaussian primitives instead of projecting them onto the image plane. By combining a $k$-buffer hit-based marching scheme with the hardware-accelerated NVIDIA OptiX interface, this approach achieves both the accuracy required for realistic camera-effect rendering and the efficiency necessary for large-scale data generation.
By integrating the dynamic scene model with differentiable ray tracing, we reconstruct dynamic scenes end-to-end via training with a color loss. At each training iteration, we randomly sample a view $v\in\mathcal{V}$ and a time $t\in\mathcal{T}$, obtain the deformed Gaussians $G_{t}$, and render them via ray tracing from view $v$ to obtain the image $C_{v,t}$. We then compute the L1 loss between $C_{v,t}$ and the ground-truth image $\hat{C}_{v,t}$, denoted by $\mathcal{L}_{1}$. Additionally, we apply a grid-based total variation loss $\mathcal{L}_{\mathrm{TV}}$ following [77, 10]. The overall loss function is defined as
@@ 3.3 Camera Effects Rendering
Fisheye. Fisheye lenses enable extremely wide-angle capture but introduce strong radial distortion. To simulate this effect, we follow Blender [6] and adopt a 4th-degree polynomial radial distortion model, which provides a flexible and physically grounded representation of fisheye lens characteristics. The model is defined as:
@@
Depth of Field (DoF). DoF controls which parts of a scene appear sharp and which appear blurred, allowing cameras to emphasize subjects at specific depths while de-emphasizing background or foreground regions. We simulate this effect using the model introduced in [14]. For each pixel, given the focus distance $f_{z}$ and the aperture radius $r_{a}$, we first compute the intersection point $\mathbf{p}=\mathbf{o}+f_{z}\mathbf{d}$ of the ideal pinhole ray with the focal plane at distance $f_{z}$. We then jitter the ray origin over a circular aperture of radius $r_{a}$ by sampling a point $\bm{\ell}=(\ell_{x},\ell_{y})$ uniformly from the unit disk and mapping it to world space using the camera’s lens-plane basis $(\mathbf{u},\mathbf{v})$, where $\mathbf{u}$ and $\mathbf{v}$ are orthogonal unit vectors spanning the lens plane. The perturbed ray origin and direction are then computed as
@@
Rolling Shutter. The rolling shutter effect arises because image rows are not captured simultaneously but sequentially over time. As a result, objects in motion or a moving camera can cause geometric distortions in the recorded image. To simulate this effect, we cast a single ray for each pixel through the deformed 3D Gaussians at its corresponding capture time. This approach resembles the motion blur technique of [14], but it differs in that we assume no motion blur: instead of integrating multiple rays over the shutter interval of a row, we trace only one ray per pixel at the corresponding sensing time. Concretely, for a pixel at row $r$, we first compute the sensing time $t_{r}$, then evaluate the deformed 3D Gaussians $G_{t_{r}}$. A ray is traced through the deformed 3D Gaussians $G_{t_{r}}$ to determine the pixel’s radiance. We currently assume a static camera; however, the method can be extended to a moving camera by interpolating the camera trajectory at time $t_{r}$ and casting the ray from the position.